OpenStudy (anonymous):

Medal for help :) The ordered pairs (1,1) (2,8) (3,27) (4,64) and (5, 125) represent which function?

OpenStudy (anonymous):

yes, because each x-value corresponds to ONLY ONE y-value

OpenStudy (dbzfan836):

he didnt say if they represented, he said what do they represent

OpenStudy (anonymous):

B

OpenStudy (dbzfan836):

D, acctually. 3*4=12, not 27

OpenStudy (dbzfan836):

2^3=8;3^3=27

OpenStudy (dbzfan836):

4^3=64

OpenStudy (dbzfan836):

anything else?

OpenStudy (anonymous):

d

OpenStudy (dbzfan836):

yes

OpenStudy (anonymous):

HAHA

OpenStudy (dbzfan836):

A fifth poem By Dbzfan836 My life is a struggle, One to be heard, Ya know why? ‘Cause why not Zoidberg? I’m stuck on the streets, Where nobody meets, And Where I keep: ‘Just why not Zoidberg?’ Then there was a day, That memes were made, So then I say, ‘Why not Zoidberg?’

OpenStudy (dbzfan836):

I made 5 other ones, this one dedicated to my favorite meme

OpenStudy (dbzfan836):

A third poem By Dbzfan836 I lay in bed; I try to sleep; I just can’t; I keep thinkin about me; Think about my day; Think about my night; Think about a nap; Think about Minecraft; Get outta bed; Get my candle read; Get to the Laptop; Play like a boss wop; Play all night; Play all day; Just for fun and Games.

OpenStudy (anonymous):

Nice good writer

OpenStudy (dbzfan836):

it takes me 5 minutes to write one, but like 10 to first think of it

OpenStudy (dbzfan836):

I make it together in my head, then write it all on paper. do some edit, and its good to go

OpenStudy (dbzfan836):

dedicated to my cat: A fourth poem By Dbzfan836 My life hurts, From losing my perch, Into my kitty paws, Because I break physic laws. I can only land on my feet, It makes me feel unique, Below my kitty paws, Because I break physic laws. I have sharp claws, Which help clean my jaws, Inside my kitty paws, Because I break Physic laws.

OpenStudy (dbzfan836):

wrote all of em in 1/2 hour

OpenStudy (dbzfan836):

when I was tired out of my mind: A sixth poem By Dbzfan836 I wanna play, I wanna play, I wanna play Every single day I wanna play, I wanna play, I wanna play At the Hudson Bay I wanna play, I wanna play, I wanna play Till the dawn of the day

OpenStudy (dbzfan836):

next question